A fishing traveler off the coast of Alaska was ordered by the U.S. Coast Guard to change course. They were to travel 6 miles west and then sail 12 miles south to miss a large iceberg before continuing on the original course. How many miles out of the way did the traveler travel?
Answer by ankor@dixie-net.com(22740)

A fishing traveler off the coast of Alaska was ordered by the U.S. Coast Guard to change course. They were to travel 6 miles west and then sail 12 miles south to miss a large iceberg before continuing on the original course. How many miles out of the way did the traveler travel?
:
This situation forms a right triangle:
The length of the original course is the hypotenuse
The legs 6 and 12 miles actually traveled, form the other two sides.
:
Using c^2 = a^2 + b^2; a = 6; b = 12; find c
:
c^2 = 6^2 + 12^2
c^2 = 36 + 144
c = Sqrt(180)
c = 13.4 mi is the distance of the original course to that point
:
Actual distance traveled to that point: 6 + 12 = 18 mi
:
Miles out of the way: 18 - 13.4 = 4.6 miles